A 4kg panel of a car oscillators with frequency 5hz amplitude of 25cm . If the oscillation is assumed to be S.H.M calculate




The maximum velocity of the panel




The total energy of the panel




The maximum potential and K.E energy of the panel is 2cm from its equlibrium equation

1
Expert's answer
2022-02-06T14:33:37-0500

Given:

"m=4\:\rm kg"

"f=5\rm\: Hz"

"A=0.25\:\rm m"


The maximum velocity of the panel


"v_{\max}=A\omega=2\pi fA
\\=2*3.14*5*0.25=7.85\:\rm m/s"

The total energy of the panel

"W=\frac{m\omega^2A^2}{2}=\frac{4\pi^2mf^2A^2}{2}"

"=\frac{4*3.14^2*4*5^2*0.25^2}{2}=123\:\rm J"

The maximum potential and K.E energy of the panel


"E_{k\max}=E_{p\max}=W=123\:\rm J"